# U2UData: A Large-scale Cooperative Perception Dataset for Swarm UAVs Autonomous Flight

### v1

@@ 1. Introduction
CoPerception-UAVs(17, ) 2022 - AirSim + Carla 5 Fixed DF, FF 4s 1 1 1 D, S
CoPerception-UAVs+(36, ) 2023 - AirSim + Carla 10 Fixed DF, FF 4s 1 1 2 D, S
@@
Perception is critical in the autonomous flight task of Unmanned Aerial Vehicles (UAVs) for accurate navigation and safe planning(11, ; 39, ). The recent development of deep learning brings significant breakthroughs in various perception tasks such as 3D object detec- tion(5, ), object tracking(6, ), and semantic segmentation(13, ). However, single-UAV vision systems still suffer from many real-world challenges, such as occlusion and short-range perception capability(17, ; 36, ; 37, ), which can cause catastrophic accidents and are key bottlenecks in improving low-altitude economic task performance. The shortcomings are mainly due to the limited field-of-view of individual UAVs, resulting in an incomplete understanding of the environment.
A growing interest and recent advances in cooperative perception systems(12, ; 14, ; 15, ; 16, ) have enabled a new paradigm that can potentially overcome the limitations of single-UAV perception. By leveraging UAV-to-UAV (U2U) technologies, multiple connected and automated UAVs (UAVs) can communicate and share captured sensor information simultaneously. As shown in Figure 1, swarm UAVs flying autonomously in a dynamic open environment, the ego UAV (blue) struggles to perceive the tracking object (deer) due to leaf obstruction and snow interference. Incorporating the multimodal features of the nearby UAVs (yellow or red) can further distribute multiple tasks and achieve greater flexibility, robustness, and perceptual range, leading to significant advantages in harsh and complex environments.
However, despite the great promise, cooperative perception is mainly focused on the vehicles and ignores the UAV literature, which remains challenging to validate U2U perception in real-world scenarios due to the lack of public datasets. Existing U2U cooperative perception datasets, as shown in Table 1, CoPerception-UAVs(17, ) and CoPerception-UAVs+(36, ) rely on open-source simulators such as AirSim(airsim, ) and CARLA(CARLA, ) and consider only 1 terrain, 1 weather, and 1 to 2 sensor types; they collect datasets using fixed altitude and consistent or fixed formation mode. In real-world scenarios, compared to autonomous driving, autonomous flight has more freedom, faces more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Obviously, there will be a clear domain gap between existing synthetic data and real-world data. As a result, models trained on these datasets may not generalize well to realistic flight situations.
@@
Our contributions can be summarized as follows:
@@ 2. Related Work
Simulator FightGear(S1, ) XPlan(S2, ) Jmavsim(S3, ) Gazebo(S4, ) AirSim(airsim, ) RflySim(S5, ) Isaac Sim(S6, ) U2USim (Our)
@@
Cooperative perception. Due to the inherent limitations of an agent’s own sensors (e.g., camera/LiDAR), occlusions, sensor degradation or failure, and long-range perception are extremely challenging for single-agent systems, with potentially disastrous consequences in harsh and complex environments. Cooperative perception systems can learn how to share information among multiple agents to perceive the environment better than individually. Existing cooperative perception approaches can be roughly divided into three categories: (1) Early Fusion(1, ). Agents transmit raw sensor data directly to other collaborators, and the ego agent makes task decisions based on the fused raw data, which preserves complete information but requires a large bandwidth. (2) Late Fusion(2, ). Each agent makes task decisions using its sensor data and delivers the decision results to others. The ego agent applies Non-maximum suppression to produce the final outputs, which maintain small transmission bandwidths but cannot achieve deep fusion of sensor information. (3) Intermediate fusion(12, ; 13, ; 14, ; 15, ; 16, ; 17, ). Adjacent UAVs utilize a neural feature extractor to derive intermediate features, which are then compressed and sent to the ego UAV for cooperative feature fusion. This method maximizes the benefits of both early and late fusion and is well-suited for extensive deployment scenarios. However, despite its great promise, cooperative perception mainly focuses on vehicles and ignores the UAV literature, which remains challenging to validate U2U perception in real-world scenarios due to the lack of public datasets.
@@
Cooperative perception datasets. Collecting datasets(19, ; 20, ; 21, ; 22, ; 23, ; 24, ; 25, ) is crucial for advancing algorithmic research. Public cooperative perception datasets have significantly accelerated progress in autonomous driving technologies in recent years. For instance, OPV2V(4, ) introduced the first 3D cooperative perception dataset, leveraging CARLA(CARLA, ) and OpenCDA(Opencda, ) joint simulation environment. DAIR-V2X pioneered real-world datasets for cooperative detection. V2x-seq(10, ) developed the initial sequential dataset for vehicle-infrastructure cooperative perception and forecasting. V2X-Sim(7, ) further explored vehicle-to-everything perception feasibility using synthesized data from the CARLA simulator(9, ). V2V4Real(8, ) established the first large-scale real-world dataset for vehicle-to-everything cooperative perception. However, cooperative perception datasets in autonomous flight scenarios remain scarce. Existing U2U cooperative perception datasets, such as CoPerception-UAVs(17, ) and CoPerception-UAVs+(36, ), rely on open-source simulators like AirSim(airsim, ) and CARLA(CARLA, ), featuring limited terrain, weather, and sensor types. These datasets collect data at fixed altitudes and in consistent or fixed formation modes. In contrast to autonomous driving, UAVs’ autonomous flight presents greater freedom, encounters more complex environments, and is more influenced by natural weather due to their smaller size. Hence, there exists a notable domain gap between existing synthetic data and real-world data, potentially limiting the generalization of models trained on these datasets to realistic flying scenarios. In this paper, we introduce the first large-scale cooperative perception dataset for swarm UAVs autonomous flight.
@@ 3. U2USim Environment
A simulator for swarm-UAVs must be able to more realistically simulate dynamic physical characteristics(S6, ) (such as collision); sensors such as IMU(S1, ), camera(S2, ; S3, ), GPS(S4, ), lidar(airsim, ) etc; and interaction with the ROS ecosystem(S5, ). In real-world scenarios, UAVs are extremely sensitive to the effects of temperature, humidity, and airflow due to their small size. Existing UAV simulators, as shown in Table 2, can visually realize digital twins(S6, ) of the real world through GPU rendering, but it is difficult to provide modal information other than visual and LiDAR modalities. Many existing studies manually set values based on experience for temperature, humidity, brightness, smoke, and airflow modalities. However, because sensors are closely related to terrain and meteorology, and have strong interactions between modalities, the model trained by existing simulators is difficult to deploy in the real world.
@@ 3.1. U2USim Design
As shown in Figure 2, U2USim uses Unreal Engine (UE) 5.2222https://www.unrealengine.com/en-US/unreal-engine-5 to construct a scaled-down 3km*3km simulated environment map based on the map of Yunnan Province. U2USim includes 4 types of terrain: mountains, hills, plains and basins. The elevation range is [56.6, 3000]m. Based on the vegetation distribution in Yunnan, 58 types of original forest vegetation assets were constructed, and more than 15 superposition methods were used to combine them, including epiphytic growth, diagonal staggered growth, and so on. Among them, the leaves of each plant will dynamically change with wind, rain, snow, and other weather conditions. U2USim takes Yunnan Province as a prototype, including 4 terrains, 7 weather conditions, and 8 sensor types. It uses the real meteorological data of Yunnan Province collected by the China Meteorological Center to map the simulation environment based on longitude and latitude.
@@
8 sensor types. The sensors in U2USim are based on the Unreal Engine, which allows cross-platform compatibility and provides accurate data due to the engine’s strong physics calculation capabilities. The LiDAR sensor primarily uses ray detection, and with the powerful engine, it can save a lot of time in tracking and calculating each ray. All image-based data, including depth and flare, are generated using Unreal rendering techniques. The flare image is created through the ID segmentation process, where objects are rendered using mapped temperature colors and output. The depth image is captured during the depth test process. The Luminance sensor uses a camera on the drone to capture live textures and calculate the average brightness and smoke concentration of each pixel. The brightness function is based on the color stimulus curve(A4, ) of the human eye. The smoke concentration function is based on the Laplacian gradient sum algorithm. To enhance the credibility of the data captured by the sensors, including temperature, humidity and airflow, weather data from Yunnan Province was used to map into the simulator environment (as shown in subsection 3.2).
@@ 4. U2Udata Dataset
CoPerception-UAVs(17, ) 2022 131.9K 800*450 - - 1.94M - - - - -
CoPerception-UAVs+(36, ) 2023 52.76K 800*450 52.76K - - - - - - -
@@
3D bounding boxes annotation. For annotating 3D bounding boxes on the gathered LiDAR data, we utilize SusTechPoint(SusTechPoint, ), a robust open-source labeling tool. We also manually refine the annotations. There are a total of three object classes, including bear, deer, and wolf. For each object, we annotate its 3D bounding box with 7 degrees of freedom, encompassing its location (x, y, z) and rotation (expressed as quaternions: w, x, y, z). The location (x, y, z) corresponds to the center of the bounding box. These 3D bounding boxes are annotated separately based on the global coordinate system of each UAV. This approach enables the sensor data from each UAV to be treated independently as a single-agent detection task. We initialize the relative pose of the two UAVs for each frame using positional information provided by the GPS on both UAVs.
@@
When2Com(16, ) 0.12 47.83 65.69 37.24 29.54 37.37 34.84 16.71 9.70
DiscoNet(15, ) 0.13 53.65 70.75 44.06 33.74 52.36 43.18 19.53 12.10
V2VNet(12, ) 0.13 59.00 71.32 47.68 35.75 48.75 36.22 22.90 13.64
V2X-ViT(14, ) 0.12 56.15 75.58 43.99 35.73 50.17 46.95 19.19 9.22
CoBEVT(13, ) 0.13 60.37 76.85 44.81 41.59 52.58 43.08 22.44 9.95
Where2com(17, ) 0.13 60.77 77.87 46.71 42.03 54.71 45.74 23.75 11.85
@@
When2Com(16, ) 0.12 46.61 66.59 30.68 17.85 16.92 30.35 10.86 0.53
DiscoNet(15, ) 0.13 52.50 72.84 40.32 19.22 21.88 39.79 16.43 0.77
V2VNet(12, ) 0.13 50.05 71.80 40.34 21.96 21.22 34.54 20.11 4.53
V2X-ViT(14, ) 0.12 49.47 75.85 35.37 18.03 24.51 44.67 14.82 3.36
CoBEVT(13, ) 0.13 50.62 76.31 36.73 22.69 22.30 40.56 17.21 5.37
Where2com(17, ) 0.13 56.45 74.84 40.45 23.94 20.28 41.83 17.42 4.26
@@ 5.1. Cooperative 3D Object Detection
Several prior works have demonstrated the efficacy of cooperative perception utilizing LiDAR sensors(17, ; 36, ; 37, ). Nevertheless, whether, when, and how this U2U cooperation using other modalities perception systems has few explored works. In this paper, each UAV first extracts LiDAR modal features through feature encoders and transmits them to adjacent UAVs. In the future, multimodal collaborative perception algorithms can be designed based on this dataset.
@@
  • Bandwidth limitation. Conventional U2U communication technologies often operate within narrow bandwidth cons- traints(12, ; 13, ; 14, ) and cooperative detection algorithms need to balance between precision and the bandwidth cost carefully.
  • Location error. Due to environmental disturbances, inherent errors(32, ) in the adjacent UAVs’ relative poses will lead to coordinate system mapping errors.
  • Asynchronicity. Heterogeneous UAVs(36, ) are inevitable and the distance between UAVs is variable. Different sensor sampling frequencies and transmission latencies will cause UAVs asynchronous.
@@
  • Early Fusion. Each UAV will transmit raw multimodal information directly to other collaborators(1, ), and the ego UAV aggregates all multimodal features to its own feature map.
  • Late Fusion. Each UAV makes task decisions utilizing its multimodal information and delivers the decision results to others(2, ), and the ego UAV applies non-maximum suppression to produce the final results.
  • Intermediate Fusion. Adjacent UAVs employ a neural feature extractor to derive intermediate features, which are subsequently compressed and broadcasted to the ego UAV for cooperative feature fusion. It can fully exploit the advantages of early and late fusion and is suitable for large-scale deployment. We benchmark some outstanding intermediate fusion methods, including When2Com(16, ), DiscoNet(15, ), V2VNet(12, ), V2X-ViT(14, ), CoBEVT(13, ), and Where2com(17, ).
Evaluation. The evaluation area extends by [-100, 100]m in both the x and y directions relative to the ego UAV. Our evaluation metric for UAV detection performance is the Average Precision (AP) at Intersection over Union (IoU) thresholds of 0.5 and 0.7. To assess transmission efficiency, we utilize the average bandwidth, which quantifies the transmitted data size as specified by the algorithm. Consistent with (36, ; ef, ), we assess all models under two conditions: 1) Synchronous setting: All UAV feature fusion modules do not consider transmission delay and UAV heterogeneity. 2) Asynchronous setting: All UAV data transmission delays are randomly selected between [0, 150]ms and are interfered by the sampling frequencies of heterogeneous UAV sensors.

### v3

@@ 1. Introduction
CoPerception-UAVs(Hu et al., 2022) 2022 - AirSim + Carla 5 Fixed DF, FF 4s 1 1 1 D, S
CoPerception-UAVs+(Hu et al., 2023) 2023 - AirSim + Carla 10 Fixed DF, FF 4s 1 1 2 D, S
@@
Perception is critical in the autonomous flight task of Unmanned Aerial Vehicles (UAVs) for accurate navigation and safe planning(Wei et al., 2024; Zhao et al., 2023). The recent development of deep learning brings significant breakthroughs in various perception tasks such as 3D object detec- tion(Zhang et al., 2022), object tracking(Sun et al., 2023), and semantic segmentation(Xu et al., 2022a). However, single-UAV vision systems still suffer from many real-world challenges, such as occlusion and short-range perception capability(Hu et al., 2023; Lu et al., 2024a), which can cause catastrophic accidents and are key bottlenecks in improving low-altitude economic task performance. The shortcomings are mainly due to the limited field-of-view of individual UAVs, resulting in an incomplete understanding of the environment.
A growing interest and recent advances in cooperative perception systems(Wang et al., 2020; Xu et al., 2022b; Li et al., 2021; Liu et al., 2020) have enabled a new paradigm that can potentially overcome the limitations of single-UAV perception. By leveraging UAV-to-UAV (U2U) technologies, multiple connected and automated UAVs (UAVs) can communicate and share captured sensor information simultaneously. As shown in Figure 1, swarm UAVs flying autonomously in a dynamic open environment, the ego UAV (blue) struggles to perceive the tracking object (deer) due to leaf obstruction and snow interference. Incorporating the multimodal features of the nearby UAVs (yellow or red) can further distribute multiple tasks and achieve greater flexibility, robustness, and perceptual range, leading to significant advantages in harsh and complex environments.
However, despite the great promise, cooperative perception is mainly focused on the vehicles and ignores the UAV literature, which remains challenging to validate U2U perception in real-world scenarios due to the lack of public datasets. Existing U2U cooperative perception datasets, as shown in Table 1, CoPerception-UAVs(Hu et al., 2022) and CoPerception-UAVs+(Hu et al., 2023) rely on open-source simulators such as AirSim(Shah et al., 2018) and CARLA(Dosovitskiy et al., 2017) and consider only 1 terrain, 1 weather, and 1 to 2 sensor types; they collect datasets using fixed altitude and consistent or fixed formation mode. In real-world scenarios, compared to autonomous driving, autonomous flight has more freedom, faces more complex environments, and is more susceptible to the influence of temperature, humidity, and airflow due to its smaller size. Obviously, there will be a clear domain gap between existing synthetic data and real-world data. As a result, models trained on these datasets may not generalize well to realistic flight situations.
@@
Our contributions [^2] https://github.com/fengtt42/U2UData can be summarized as follows:
@@ 2. Related Work
Simulator FightGear(DigitalOcean, [n. d.]) XPlan(Research, [n. d.]) Jmavsim(PX4, [n. d.]) Gazebo(Robotics, [n. d.]) AirSim(Shah et al., 2018) RflySim(FEISILAB, [n. d.]) Isaac Sim(NVIDIA, [n. d.]) U2USim (Our)
@@
Cooperative perception. Due to the inherent limitations of an agent’s own sensors (e.g., camera/LiDAR), occlusions, sensor degradation or failure, and long-range perception are extremely challenging for single-agent systems, with potentially disastrous consequences in harsh and complex environments. Cooperative perception systems can learn how to share information among multiple agents to perceive the environment better than individually. Existing cooperative perception approaches can be roughly divided into three categories: (1) Early Fusion(Chen et al., 2019). Agents transmit raw sensor data directly to other collaborators, and the ego agent makes task decisions based on the fused raw data, which preserves complete information but requires a large bandwidth. (2) Late Fusion(Rawashdeh and Wang, 2018). Each agent makes task decisions using its sensor data and delivers the decision results to others. The ego agent applies Non-maximum suppression to produce the final outputs, which maintain small transmission bandwidths but cannot achieve deep fusion of sensor information. (3) Intermediate fusion(Wang et al., 2020; Xu et al., 2022a, b; Li et al., 2021; Liu et al., 2020; Hu et al., 2022). Adjacent UAVs utilize a neural feature extractor to derive intermediate features, which are then compressed and sent to the ego UAV for cooperative feature fusion. This method maximizes the benefits of both early and late fusion and is well-suited for extensive deployment scenarios. However, despite its great promise, cooperative perception mainly focuses on vehicles and ignores the UAV literature, which remains challenging to validate U2U perception in real-world scenarios due to the lack of public datasets.
@@
Cooperative perception datasets. Collecting datasets(Lin et al., 2023; Chen et al., 2023b; Zhu et al., 2023; Ying et al., 2023; Chen et al., 2023a; Jin et al., 2023; Liu et al., 2023) is crucial for advancing algorithmic research. Public cooperative perception datasets have significantly accelerated progress in autonomous driving technologies in recent years. For instance, OPV2V(Xu et al., 2022c) introduced the first 3D cooperative perception dataset, leveraging CARLA(Dosovitskiy et al., 2017) and OpenCDA(Xu et al., 2021) joint simulation environment. DAIR-V2X pioneered real-world datasets for cooperative detection. V2x-seq(Yu et al., 2023) developed the initial sequential dataset for vehicle-infrastructure cooperative perception and forecasting. V2X-Sim(Li et al., 2022) further explored vehicle-to-everything perception feasibility using synthesized data from the CARLA simulator(Yu et al., 2022). V2V4Real(Xu et al., 2023) established the first large-scale real-world dataset for vehicle-to-everything cooperative perception. However, cooperative perception datasets in autonomous flight scenarios remain scarce. Existing U2U cooperative perception datasets, such as CoPerception-UAVs(Hu et al., 2022) and CoPerception-UAVs+(Hu et al., 2023), rely on open-source simulators like AirSim(Shah et al., 2018) and CARLA(Dosovitskiy et al., 2017), featuring limited terrain, weather, and sensor types. These datasets collect data at fixed altitudes and in consistent or fixed formation modes. In contrast to autonomous driving, UAVs’ autonomous flight presents greater freedom, encounters more complex environments, and is more influenced by natural weather due to their smaller size. Hence, there exists a notable domain gap between existing synthetic data and real-world data, potentially limiting the generalization of models trained. In this paper, we introduce the first large-scale cooperative perception dataset for swarm UAVs autonomous flight.
@@ 3. U2USim Environment
A simulator for swarm-UAVs must be able to more realistically simulate dynamic physical characteristics(NVIDIA, [n. d.]) (such as collision); sensors such as IMU(DigitalOcean, [n. d.]), camera(Research, [n. d.]; PX4, [n. d.]), GPS(Robotics, [n. d.]), lidar(Shah et al., 2018) etc; and interaction with the ROS ecosystem(FEISILAB, [n. d.]). In real-world scenarios, UAVs are extremely sensitive to the effects of temperature, humidity, and airflow due to their small size. Existing UAV simulators, as shown in Table 2, can visually realize digital twins(NVIDIA, [n. d.]) of the real world through GPU rendering, but it is difficult to provide modal information other than visual and LiDAR modalities. Many existing studies manually set values based on experience for temperature, humidity, brightness, smoke, and airflow modalities. However, because sensors are closely related to terrain and meteorology, and have strong interactions between modalities, the model trained by existing simulators is difficult to deploy in the real world.
@@ 3.1. U2USim Design
As shown in Figure 2, U2USim uses Unreal Engine (UE) 5.2333https://www.unrealengine.com/en-US/unreal-engine-5 to construct a scaled-down 3km*3km simulated environment map based on the map of Yunnan Province. U2USim includes 4 types of terrain: mountains, hills, plains and basins. The elevation range is [56.6, 3000]m. Based on the vegetation distribution in Yunnan, 58 types of original forest vegetation assets were constructed, and more than 15 superposition methods were used to combine them, including epiphytic growth, diagonal staggered growth, and so on. Among them, the leaves of each plant will dynamically change with wind, rain, snow, and other weather conditions. U2USim takes Yunnan Province as a prototype, including 4 terrains, 7 weather conditions, and 8 sensor types. It uses the real meteorological data of Yunnan Province collected by the China Meteorological Center to map the simulation environment based on longitude and latitude.
@@
8 sensor types. The sensors in U2USim are based on the Unreal Engine, which allows cross-platform compatibility and provides accurate data due to the engine’s strong physics calculation capabilities. The LiDAR sensor primarily uses ray detection, and with the powerful engine, it can save a lot of time in tracking and calculating each ray. All image-based data, including depth and flare, are generated using Unreal rendering techniques. The flare image is created through the ID segmentation process, where objects are rendered using mapped temperature colors and output. The depth image is captured during the depth test process. The Luminance sensor uses a camera on the drone to capture live textures and calculate the average brightness and smoke concentration of each pixel. The brightness function is based on the color stimulus curve(Liao et al., 2022) of the human eye. The smoke concentration function is based on the Laplacian gradient sum algorithm. To enhance the credibility of the data captured by the sensors, including temperature, humidity and airflow, weather data from Yunnan Province was used to map into the simulator environment (as shown in subsection 3.2).
@@ 4. U2Udata Dataset
CoPerception-UAVs(Hu et al., 2022) 2022 131.9K 800*450 - - 1.94M - - - - -
CoPerception-UAVs+(Hu et al., 2023) 2023 52.76K 800*450 52.76K - - - - - - -
@@
3D bounding boxes annotation. For annotating 3D bounding boxes on the gathered LiDAR data, we utilize SusTechPoint(Li et al., 2020), a robust open-source labeling tool. We also manually refine the annotations. There are a total of three object classes, including bear, deer, and wolf. For each object, we annotate its 3D bounding box with 7 degrees of freedom, encompassing its location (x, y, z) and rotation (expressed as quaternions: w, x, y, z). The location (x, y, z) corresponds to the center of the bounding box. These 3D bounding boxes are annotated separately based on the global coordinate system of each UAV. This approach enables the sensor data from each UAV to be treated independently as a single-agent detection task. We initialize the relative pose of the two UAVs for each frame using positional information provided by the GPS on both UAVs.
@@
When2Com(Liu et al., 2020) 0.12 47.83 65.69 37.24 29.54 37.37 34.84 16.71 9.70
DiscoNet(Li et al., 2021) 0.13 53.65 70.75 44.06 33.74 52.36 43.18 19.53 12.10
V2VNet(Wang et al., 2020) 0.13 59.00 71.32 47.68 35.75 48.75 36.22 22.90 13.64
V2X-ViT(Xu et al., 2022b) 0.12 56.15 75.58 43.99 35.73 50.17 46.95 19.19 9.22
CoBEVT(Xu et al., 2022a) 0.13 60.37 76.85 44.81 41.59 52.58 43.08 22.44 9.95
Where2com(Hu et al., 2022) 0.13 60.77 77.87 46.71 42.03 54.71 45.74 23.75 11.85
@@
When2Com(Liu et al., 2020) 0.12 46.61 66.59 30.68 17.85 16.92 30.35 10.86 0.53
DiscoNet(Li et al., 2021) 0.13 52.50 72.84 40.32 19.22 21.88 39.79 16.43 0.77
V2VNet(Wang et al., 2020) 0.13 50.05 71.80 40.34 21.96 21.22 34.54 20.11 4.53
V2X-ViT(Xu et al., 2022b) 0.12 49.47 75.85 35.37 18.03 24.51 44.67 14.82 3.36
CoBEVT(Xu et al., 2022a) 0.13 50.62 76.31 36.73 22.69 22.30 40.56 17.21 5.37
Where2com(Hu et al., 2022) 0.13 56.45 74.84 40.45 23.94 20.28 41.83 17.42 4.26
@@ 5.1. Cooperative 3D Object Detection
Several prior works have demonstrated the efficacy of cooperative perception utilizing LiDAR sensors(Hu et al., 2022, 2023; Lu et al., 2024a). Nevertheless, whether, when, and how this U2U cooperation using other modalities perception systems has few explored works. In this paper, each UAV first extracts LiDAR modal features through feature encoders and transmits them to adjacent UAVs. In the future, multimodal collaborative perception algorithms can be designed based on this dataset.
@@
  • Bandwidth limitation. Conventional U2U communication technologies often operate within narrow bandwidth cons- traints(Wang et al., 2020; Xu et al., 2022a) and cooperative detection algorithms need to balance between precision and the bandwidth cost carefully.
  • Location error. Due to environmental disturbances, inherent errors(Wu et al., 2024) in the adjacent UAVs’ relative poses will lead to coordinate system mapping errors.
  • Asynchronicity. Heterogeneous UAVs(Hu et al., 2023) are inevitable and the distance between UAVs is variable. Different sensor sampling frequencies and transmission latencies will cause UAVs asynchronous.
@@
  • Early Fusion. Each UAV will transmit raw multimodal information directly to other collaborators(Chen et al., 2019), and the ego UAV aggregates all multimodal features to its own feature map.
  • Late Fusion. Each UAV makes task decisions utilizing its multimodal information and delivers the decision results to others(Rawashdeh and Wang, 2018), and the ego UAV applies non-maximum suppression to produce the final results.
  • Intermediate Fusion. Adjacent UAVs employ a neural feature extractor to derive intermediate features, which are subsequently compressed and broadcasted to the ego UAV for cooperative feature fusion. It can fully exploit the advantages of early and late fusion and is suitable for large-scale deployment. We benchmark some outstanding intermediate fusion methods, including When2Com(Liu et al., 2020), DiscoNet(Li et al., 2021), V2VNet(Wang et al., 2020), V2X-ViT(Xu et al., 2022b), CoBEVT(Xu et al., 2022a), and Where2com(Hu et al., 2022).
Evaluation. The evaluation area extends by [-100, 100]m in both the x and y directions relative to the ego UAV. Our evaluation metric for UAV detection performance is the Average Precision (AP) at Intersection over Union (IoU) thresholds of 0.5 and 0.7. To assess transmission efficiency, we utilize the average bandwidth, which quantifies the transmitted data size as specified by the algorithm. Consistent with (Hu et al., 2023; Lu et al., 2024b), we assess all models under two conditions: 1) Synchronous setting: All UAV feature fusion modules do not consider transmission delay and UAV heterogeneity. 2) Asynchronous setting: All UAV data transmission delays are randomly selected between [0, 150]ms and are interfered by the sampling frequencies of heterogeneous UAV sensors.